# The Type Ia Supernova 2021hem: A 2003fg-like Event in an Apparently Hostless Environment

### v1

@@
11institutetext: Department of Physics and Astronomy, Aarhus University, Ny Munkegade 120, DK-8000 Aarhus C, Denmark (11email: email@subhashbose.com, max@phys.au.dk) 22institutetext: Department of Astronomy, University of Michigan, Ann Arbor, MI 48109, USA 33institutetext: INAF – Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy 44institutetext: Cosmic DAWN Center, Niels Bohr Institute, University of Copenhagen, Lyngbyvej 2, 2100 Copenhagen Ø, Denmark 55institutetext: Institute for Astronomy, University of Hawai’i, 2680 Woodlawn Drive, Honolulu HI 96822, USA 66institutetext: Observatories of the Carnegie Institution for Science, 813 Santa Barbara St, Pasadena, CA, 91101, USA 77institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218-2410, USA 88institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, E-08193 Barcelona, Spain 99institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), E-08034 Barcelona, Spain 1010institutetext: Department of Physics, Chalmers University of Technology, SE-412 96, Göteborg, Sweden 1111institutetext: European Southern Observatory, Alonso de Córdova 3107, Casilla 19, Santiago, Chile 1212institutetext: Planetary Science Institute, 1700 East Fort Lowell Road, Suite 106, Tucson, AZ 85719-2395, USA 1313institutetext: CENTRA, Instituto Superior Técnico, Universidade de Lisboa, Av. Rovisco Pais 1, 1049-001 Lisboa, Portugal 1414institutetext: Astronomical Observatory, University of Warsaw, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 1515institutetext: Cardiff Hub for Astrophysics Research and Technology, School of Physics & Astronomy, Cardiff University, Queens Buildings, The Parade, Cardiff, CF24 3AA, UK 1616institutetext: Astrophysics Research Institute, Liverpool John Moores University, IC2, Liverpool Science Park, 146 Brownlow Hill, Liverpool L3 5RF, UK 1717institutetext: School of Physics, Trinity College Dublin, The University of Dublin, Dublin 2, Ireland 1818institutetext: Instituto de Ciencias Exactas y Naturales (ICEN), Universidad Arturo Prat, Chile 1919institutetext: The Oskar Klein Centre, Department of Physics, Stockholm University, AlbaNova, 10691 Stockholm, Sweden 2020institutetext: INAF - Osservatorio Astronomico di Capodimonte, Salita Moiariello 16, 80131 Napoli, Italy 2121institutetext: Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721-0065, USA 2222institutetext: Finnish Centre for Astronomy with ESO (FINCA), Quantum, Vesilinnantie 5, University of Turku, 20014 Turku, Finland 2323institutetext: Department of Physics and Astronomy, University of Turku, 20014 Turku, Finland
@@
Subhash Bose[Uncaptioned image]    M. D. Stritzinger[Uncaptioned image]    A. Malmgaard    C. J. Miller[Uncaptioned image]    N. Elias-Rosa[Uncaptioned image]    J. P. U. Fynbo[Uncaptioned image]    C. Ashall[Uncaptioned image]    C. R. Burns[Uncaptioned image]    J. M. DerKacy[Uncaptioned image]    L. Galbany[Uncaptioned image]    C. P. Gutiérrez[Uncaptioned image]    W. B. Hoogendam[Uncaptioned image]    E. A. M. Jensen[Uncaptioned image]    K. Medler[Uncaptioned image]    Alaa Alburai[Uncaptioned image]    J. Anderson[Uncaptioned image]    E. Baron[Uncaptioned image]    J. Duarte[Uncaptioned image]    M. Gromadzki[Uncaptioned image]    C. Inserra[Uncaptioned image]    P. A. Mazzali[Uncaptioned image]    T. E. Müller-Bravo[Uncaptioned image]    P. Lundqvist    A. Reguitti[Uncaptioned image]    I. Salmaso[Uncaptioned image]    D. J. Sand[Uncaptioned image]    G. Valerin[Uncaptioned image]
(Received November 10, 2025 Accepted $X$ Month, 2026.)
@@
We report observations of a Type Ia supernova 2021hem, discovered within 48 hours of last non-detection, and located in an apparently hostless environment. With a peak absolute $B$-band magnitude of $M_{B,\max}=-19.96\pm 0.29$ mag, SN 2021hem lies at the luminous end of the SN Ia distribution. Its near-infrared and $i$-band light curves lack the secondary maximum, which is otherwise ubiquitous to normal and 1991T-like SNe Ia. Instead, these properties make SN 2021hem closely resemble 2003fg-like events. The slowly evolving light curves (characterized by $\Delta m_{15}(B)=1.02\pm 0.02$ mag; $s_{BV}=0.94\pm 0.05$), and the earliest spectrum showing C ii $\lambda$6580 and $\lambda$7235 absorption lines, further support this classification. Other spectroscopic features, including Si ii line diagnostics, resemble normal SNe Ia. A fit of a “fireball” model to the early-time light curves yields a time of first light of $t_{\rm first}=-16.43^{+0.45}_{-0.38}$ days relative to $B$-band maximum. The first photometric detection occurs $1.51^{+0.45}_{-0.38}$ days before the onset of fireball-like flux rise. This early emission, together with the intrinsic $(g-r)_{0}$ color, is inconsistent with circumstellar or companion interaction. Instead, shallow 56Ni mixing or an asymmetric 56Ni distribution offers a plausible explanation for the delayed onset of the fireball flux rise, while a double-detonation scenario with a thin helium shell remains a less likely alternative. Notably, SN 2021hem represents the fifth known 2003fg-like SN that has early-time activity or excess flux emission. The estimated mass of radioactive 56Ni synthesized in SN 2021hem is $1.00\pm 0.09$ M. Deep GTC imaging obtained 2.5 years after the explosion, with an estimated limiting magnitude of $m_{lim,r}=24.4$ mag and a surface-brightness limit of $\mu_{lim,r}=26.3$$\mathrm{mag\penalty 10000\ arcsec^{-2}}$, reveals no coincident host, thereby ruling out most faint dwarf and ultra-diffuse galaxies (UDGs). Alternatively, if the nearest plausible host galaxy located at a projected distance of 102  kpc is assumed, the progenitor would need to be a hyper-velocity star ejected at $\sim 2200$$\rm{\,km\,s^{-1}}$ from the host. A faint diffuse feature $\approx 6$ kpc from the SN site has also been detected in the GTC image, with its surface brightness within the limits of UDGs. However, it is unclear whether it is a galaxy and is associated with SN 2021hem. Considering its large normalized directional light distance ($d_{\mathrm{DLR}}$ $\sim 3-4$) from the SN, and its unusual elongation, it is a candidate of low probability to be the host galaxy of SN 2021hem. These results identify SN 2021hem as one of the strongest candidates for a hostless SN Ia, underscoring the diversity of luminous, slowly evolving, 2003fg-like explosions and the wide range of environments in which they may occur.
@@ 2 Discovery and classification
SN 2021hem was identified by ALeRCE (Automatic Learning for the Rapid Classification of Events) broker system with J2000 coordinates R.A. $16^{h}21^{m}16^{s}.01$ and Decl. $14^{\circ}33^{m}09^{s}.61$ (Forster2021). This discovery was made possible with ALeRCE’s stamp classifier (Carrasco-Davis2021), utilizing data from the Zwicky Transient Facility (ZTF; Bellm2019) public data stream. According to the Transient Name Server (TNS) discovery certificate, the first detection of the object appeared in an $r_{ZTF}$-band image taken on 2021-03-24T10:05:20.00 UT and the last non-detection was obtained 45 hours earlier (i.e, 2021-03-22T11:33:32 UT). Making use of the ZTF forced photometry service (Masci2023), we recomputed photometry of the ZTF data stream which indicates at the time of discovery of SN 2021hem had an $r_{ztf}$-band magnitude of 19.9, while the limiting magnitude of the non-detection image was 20.12 mag.
@@
Figure 1: An $r$-band image of SN 2021hem at peak brightness captured with the Nordic Optical Telescope. The location of SN 2021hem is positioned at the center in both panels. Two of the closest cataloged galaxies are marked in blue, both at the projected distance of roughly $\sim 100\mbox{\penalty 10000\ kpc}$ (see Sect. 5.1)
@@
Figure 2: The large scale structure around the position of SN 2021hem (marked with a cyan cross) shown in a field of a size of $\times2$ square degrees DSS image. The colored circles indicate the galaxies with measured redshifts from SIMBAD within redhsift range of 0.02 to 0.05. The color-bar indicates the redshifts for each color. There appears to be an overdensity of galaxies with redshifts close to $z\sim 0.027$, $\sim 0.035$ within a radius of 1 Mpc from the position of the SN. The open star symbol corresponds to the elliptical galaxy LEDA 1461683, which has the shortest projected distance of 102 kpc from the SN.
@@ 3.3 Reddening
The Milky Way reddening color excess along the line of sight toward SN 2021hem is $E(B-V)_{MW}=0.0501\pm 0.0034$ mag (2011ApJ...737..103S). The post-peak $(B-V)_{0}$ color curve of SN 2021hem, after correcting for Milky way reddening is consistent with the intrinsic color-curve of the archetypal normal SN 2011fe as well as the subluminous SN 1991bg (see Fig. 12), suggesting SN 2021hem suffered minimal host extinction. Further corroborating this, SN 2021hem spectra reveal no Na i D absorption at the redshift of the SN. The minimal host reddening is also consistent with the apparent hostless environment of the SN. Therefore, we adopt the previously stated Milky Way reddening value in our extinction correction for SN 2021hem. Assuming $R_{V}=3.1$, this corresponds to an $A^{tot}_{V}=0.155\pm 0.011$ mag.
@@ 4.1.1 Light-curve properties and template fitting
Using spline fits, the estimated peak apparent magnitude is $m_{B,max}=16.34\pm 0.01$ mag, and the post-peak light curve decline is estimated to be $\mbox{$\Delta m_{15}(B)$}=1.02\pm 0.02$ mag. After correcting for Milky Way reddening, computed K-correction of 0.005-0.005 mag, and luminosity distance for the adopted redshift of $z=0.0363\pm 0.0044$, the peak absolute magnitude of SN 2021hem is $M_{B,max}=-19.96\pm 0.29$. On constructing $(B-V)_{0}$ intrinsic color-curve with interpolated spline fits after correcting for reddening, the color-curve peaks at $28.3\pm 1.1$ day, which corresponds to color-stretch parameter (2014ApJ...789...32B) $\mbox{$s_{BV}$}=0.94\pm 0.05$.
@@
The ‘max_model’ provides fit parameters of the time of maximum as ${\rm JD}\penalty 10000\ 2459316.74\pm 0.36$ and a color-stretch parameter $s_{BV}=1.19\pm 0.04$. Table 2 provides results from direct GP splines to the optical light curves, and the estimated peak apparent magnitudes of the optical/NIR light curves estimated from the SNooPymax_model’ fits, which include corrections for Milky Way reddening and computed K-correction. Adopting a redshift of $z=0.0363\pm 0.0044$ for SN 2021hem (see Sect. 3.4), the peak apparent magnitudes computed by the ‘max_model’ indicate that it reached an absolute $B$-band peak magnitude of $M_{B}=-19.90\pm 0.29$ mag, roughly half a magnitude brighter than typical SNe Ia of similar color-stretch values.
@@
Filter JD peak Peak app. Mag. $\Delta m_{15}$
@@ 4.1.2 Estimating the time of first light, $t_{first}$
This pre-fireball emission should not be confused with the early flux excess seen in many early SNe Ia light curves, as the two phenomena are characteristically distinct. The early flux excess is typically attributed to the interaction of the SN ejecta with a non-degenerate companion star (2010ApJ...708.1025K), or with circumstellar material (e.g., Piro2016; Maeda2023; 2023MNRAS.522.6035M; 2018ApJ...864L..35S). It is estimated that approximately $20-30\%$ of all SNe Ia exhibit such early-excess features (2022MNRAS.513.3035M; 2022MNRAS.512.1317D). However, in certain subclasses of SNe Ia – such as the peculiar 2002es-like, 2003fg-like, and luminous 1991T-like events – the occurrence rate of early-excess emission appears to be considerably higher (2018ApJ...865..149J; 2024ApJ...966..139H). Crucially, unlike these early-excess signatures, the pre-fireball emission observed in SN 2021hem likely arises from a distinct physical mechanism, as discussed earlier. An early flux excess due to interaction typically emerges within a few days of explosion and appears superimposed on the power-law rise of the early light curve. This is because the short-lived excess flux from the interaction and the underlying “fireball” emission are both consequences of the expanding ejecta. In other words, the interaction can not occur without the “fireball” expansion. In contrast, the early emission seen in SN 2021hem occurs even before the onset of the fireball’s power-law flux rise, clearly distinguishing it from the early excesses or bumps arising from interaction in other SNe Ia. Moreover, 2017Natur.550...80J argued, based on their models, that an early excess arising from interaction would inevitably show a blue color evolution, as was observed in the $(\mbox{$B\!-\!V$})$ color of SN iPTF14atg. In the case of SN 2021hem the $(g-r)_{0}$ color of $0.16\pm 0.33$ mag, computed from first epoch of $g$ and $r$-band light curves, does not show any blue excess. This color is broadly consistent with, or perhaps marginally redder than, the early $(g-r)_{0}$ colors typically observed in normal SNe Ia (2020ApJ...902...48B). Therefore, considering all these factors, the interaction scenario is disfavored for the pre-fireball emission seen in SN 2021hem.
@@
Regardless of the origin, this early activity, which can be described as early flux excess or pre-fireball flux emission as in SN 2021hem, appears to be a ubiquitous feature among 2003fg-like SNe. In addition to SN 2021hem, other 2003fg-like supernovae with early-time observations and adequate photometric cadence – SNe 2020hvf, 2021zny, 2022ilv, and 2021qvo (2024ApJ...966..139H; 2025arXiv250813263A) – have all exhibited an early bump in their light curves.
@@
Figure 5: The UVOIR bolometric light curve of SN 2021hem fit with a radioactive 56Ni decay model. The time of explosion is taken as the epoch of first detection, which is $\approx 17.9$ days before the $B$-band maximum. Only data beyond 60 days are used for fitting the model. The UVOIR light curve of SN 2021hem (top panel), its time-weighted integrated luminosity (middle panel), and the ratio $t^{2}L/\left(\int Lt\penalty 10000\ dt\right)$ (bottom panel), which is defined independent of the 56Ni mass are shown.

### v2

@@
11institutetext: Department of Physics and Astronomy, Aarhus University, Ny Munkegade 120, DK-8000 Aarhus C, Denmark (11email: email@subhashbose.com, max@phys.au.dk) 22institutetext: Department of Astronomy, University of Michigan, Ann Arbor, MI 48109, USA 33institutetext: INAF – Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy 44institutetext: Cosmic DAWN Center, Niels Bohr Institute, University of Copenhagen, Lyngbyvej 2, 2100 Copenhagen Ø, Denmark 55institutetext: Institute for Astronomy, University of Hawai’i, 2680 Woodlawn Drive, Honolulu HI 96822, USA 66institutetext: Observatories of the Carnegie Institution for Science, 813 Santa Barbara St, Pasadena, CA, 91101, USA 77institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218-2410, USA 88institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, E-08193 Barcelona, Spain 99institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), E-08034 Barcelona, Spain 1010institutetext: Department of Physics, Florida State University, 77 Chieftan Way, Tallahassee, FL 32306, USA 1111institutetext: Department of Physics, Chalmers University of Technology, SE-412 96, Göteborg, Sweden 1212institutetext: European Southern Observatory, Alonso de Córdova 3107, Casilla 19, Santiago, Chile 1313institutetext: Planetary Science Institute, 1700 East Fort Lowell Road, Suite 106, Tucson, AZ 85719-2395, USA 1414institutetext: CENTRA, Instituto Superior Técnico, Universidade de Lisboa, Av. Rovisco Pais 1, 1049-001 Lisboa, Portugal 1515institutetext: Astronomical Observatory, University of Warsaw, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 1616institutetext: Cardiff Hub for Astrophysics Research and Technology, School of Physics & Astronomy, Cardiff University, Queens Buildings, The Parade, Cardiff, CF24 3AA, UK 1717institutetext: Astrophysics Research Institute, Liverpool John Moores University, IC2, Liverpool Science Park, 146 Brownlow Hill, Liverpool L3 5RF, UK 1818institutetext: School of Physics, Trinity College Dublin, The University of Dublin, Dublin 2, Ireland 1919institutetext: Instituto de Ciencias Exactas y Naturales (ICEN), Universidad Arturo Prat, Chile 2020institutetext: The Oskar Klein Centre, Department of Physics, Stockholm University, AlbaNova, 10691 Stockholm, Sweden 2121institutetext: INAF - Osservatorio Astronomico di Capodimonte, Salita Moiariello 16, 80131 Napoli, Italy 2222institutetext: Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721-0065, USA
@@
Subhash Bose[Uncaptioned image]    M. D. Stritzinger[Uncaptioned image]    A. Malmgaard    C. J. Miller[Uncaptioned image]    N. Elias-Rosa[Uncaptioned image]    J. P. U. Fynbo[Uncaptioned image]    C. Ashall[Uncaptioned image]    C. R. Burns[Uncaptioned image]    J. M. DerKacy[Uncaptioned image]    L. Galbany[Uncaptioned image]    C. P. Gutiérrez[Uncaptioned image]    W. B. Hoogendam[Uncaptioned image]    E. Y. Hsiao[Uncaptioned image]    E. A. M. Jensen[Uncaptioned image]    K. Medler[Uncaptioned image]    Alaa Alburai[Uncaptioned image]    J. Anderson[Uncaptioned image]    E. Baron[Uncaptioned image]    J. Duarte[Uncaptioned image]    M. Gromadzki[Uncaptioned image]    C. Inserra[Uncaptioned image]    P. A. Mazzali[Uncaptioned image]    T. E. Müller-Bravo[Uncaptioned image]    P. Lundqvist    A. Reguitti[Uncaptioned image]    I. Salmaso[Uncaptioned image]    D. J. Sand[Uncaptioned image]    G. Valerin[Uncaptioned image]
(Received 10 November 2025 / Accepted 24 November 2025.)
@@
We report observations of a Type Ia supernova 2021hem, discovered within 48 hours of last non-detection, and located in an apparently hostless environment. With a peak absolute $B$-band magnitude of $M_{B,\max}=-19.96\pm 0.29$ mag, SN 2021hem lies at the luminous end of the SN Ia distribution. Its near-infrared and $i$-band light curves lack the secondary maximum, which is otherwise ubiquitous to normal and 1991T-like SNe Ia. Instead, these properties make SN 2021hem closely resemble 2003fg-like events. The slowly evolving light curves (characterized by $\Delta m_{15}(B)=1.02\pm 0.02$ mag; $s_{BV}=0.94\pm 0.05$), and the earliest spectrum showing C ii $\lambda$6580 and $\lambda$7235 absorption lines, further support this classification. Other spectroscopic features, including Si ii line diagnostics, resemble normal SNe Ia. A fit of a “fireball” model to the early-time light curves yields a time of first light of $t_{\rm first}=-16.43^{+0.45}_{-0.38}$ days relative to $B$-band maximum. The first photometric detection occurs $1.51^{+0.45}_{-0.38}$ days before the onset of fireball-like flux rise. This early emission, together with the intrinsic $(g-r)_{0}$ color, is inconsistent with circumstellar or companion interaction. Instead, shallow 56Ni mixing or an asymmetric 56Ni distribution offers a plausible explanation for the delayed onset of the fireball flux rise, while a double-detonation scenario with a thin helium shell remains a less likely alternative. Notably, SN 2021hem represents the fifth known 2003fg-like SN that has early-time activity or excess flux emission. The estimated mass of radioactive 56Ni synthesized in SN 2021hem is $1.00\pm 0.09$ M. Deep GTC imaging obtained 2.5 years after the explosion, with an estimated limiting magnitude of $m_{lim,r}=24.4$ mag and a surface-brightness limit of $\mu_{lim,r}=26.3$$\mathrm{mag\penalty 10000\ arcsec^{-2}}$, reveals no coincident host, thereby ruling out most faint dwarf and ultra-diffuse galaxies (UDGs). Alternatively, if the nearest plausible AGN host galaxy located at a projected distance of 104  kpc is assumed, the progenitor would need to be a hyper-velocity star ejected at $\sim 2200$$\rm{\,km\,s^{-1}}$ from the host by AGN interaction. A faint diffuse feature $\approx 6$ kpc from the SN site has also been detected in the GTC image, with its surface brightness within the limits of UDGs. However, it is unclear whether it is a galaxy and is associated with SN 2021hem. Considering its large normalized directional light distance ($d_{\mathrm{DLR}}$ $\sim 3-4$) from the SN, and its unusual elongation, it is a candidate of low probability to be the host galaxy of SN 2021hem. These results identify SN 2021hem as one of the strongest candidates for a hostless SN Ia, underscoring the diversity of luminous, slowly evolving, 2003fg-like explosions and the wide range of environments in which they may occur.
@@ 2 Discovery and classification
SN 2021hem was identified by ALeRCE (Automatic Learning for the Rapid Classification of Events) broker system with J2000 coordinates R.A. $16^{h}21^{m}16^{s}.01$ and Decl. $14^{\circ}33^{\prime}09\aas@@fstack{\prime\prime}61$ (Forster2021). This discovery was made possible with ALeRCE’s stamp classifier (Carrasco-Davis2021), utilizing data from the Zwicky Transient Facility (ZTF; Bellm2019) public data stream. According to the Transient Name Server (TNS) discovery certificate, the first detection of the object appeared in an $r_{ZTF}$-band image taken on 2021-03-24T10:05:20.00 UT and the last non-detection was obtained 45 hours earlier (i.e, 2021-03-22T11:33:32 UT). Making use of the ZTF forced photometry service (Masci2023), we recomputed photometry of the ZTF data stream which indicates at the time of discovery of SN 2021hem had an $r_{ztf}$-band magnitude of 19.9, while the limiting magnitude of the non-detection image was 20.12 mag.
@@
Figure 1: $r$-band image of SN 2021hem at peak brightness observed with the Nordic Optical Telescope. The location of SN 2021hem is positioned at the center in both panels. Two of the closest cataloged galaxies are marked in blue, both at the projected distance of roughly $\sim 100\mbox{\penalty 10000\ kpc}$ (see Sect. 5.1)
@@
Figure 2: Large-scale structure around the position of SN 2021hem (marked with a cyan cross) shown in a field of a size of $\times2$ square degrees DSS image. The colored circles indicate the galaxies with measured redshifts from SIMBAD catalog within redshift range of 0.02 to 0.05. The color-bar indicates the redshifts for each color. There appears to be an overdensity of galaxies with redshifts close to $z\sim 0.027$, $\sim 0.035$ within a radius of 1 Mpc from the position of the SN. The open star symbol corresponds to the AGN galaxy 2MASX J16212572+1431537, which has the projected distance of 104 kpc from the SN.
@@ 3.3 Reddening
The Milky Way reddening color excess along the line-of-sight toward SN 2021hem is $E(B-V)_{MW}=0.0501\pm 0.0034$ mag (2011ApJ...737..103S). The post-peak $(B-V)_{0}$ color curve of SN 2021hem, after correcting for Milky way reddening is consistent with the intrinsic color-curve of the archetypal normal SN 2011fe as well as the subluminous SN 1991bg (see Fig. 12), suggesting SN 2021hem suffered minimal host extinction. Further corroborating this, SN 2021hem spectra reveal no Na i D absorption at the redshift of the SN. The minimal host reddening is also consistent with the apparent hostless environment of the SN. Therefore, we adopt the previously stated Milky Way reddening value in our extinction correction for SN 2021hem. Assuming $R_{V}=3.1$, this corresponds to an $A^{tot}_{V}=0.155\pm 0.011$ mag.
@@ 4.1.1 Light-curve properties and template fitting
Using spline fits, the estimated peak apparent magnitude is $m_{B,max}=16.34\pm 0.01$ mag, and the post-peak light curve decline is estimated to be $\mbox{$\Delta m_{15}(B)$}=1.02\pm 0.02$ mag. After correcting for Milky Way reddening, computed K-correction of 0.005-0.005 mag, and luminosity distance for the adopted redshift of $z=0.0363\pm 0.0049$, the peak absolute magnitude of SN 2021hem is $M_{B,max}=-19.96\pm 0.29$. On constructing $(B-V)_{0}$ intrinsic color-curve with interpolated spline fits after correcting for reddening, the color-curve peaks at $28.3\pm 1.1$ day, which corresponds to color-stretch parameter (2014ApJ...789...32B) $\mbox{$s_{BV}$}=0.94\pm 0.05$.
@@
The ‘max_model’ provides fit parameters of the time of maximum as ${\rm JD}\penalty 10000\ 2459316.74\pm 0.36$ and a color-stretch parameter $s_{BV}=1.19\pm 0.04$. Table 2 provides results from direct GP splines to the optical light curves, and the estimated peak apparent magnitudes of the optical/NIR light curves estimated from the SNooPymax_model’ fits, which include corrections for Milky Way reddening and computed K-correction. Adopting a redshift of $z=0.0363\pm 0.0049$ for SN 2021hem (see Sect. 3.4), the peak apparent magnitudes computed by the ‘max_model’ indicate that it reached an absolute $B$-band peak magnitude of $M_{B}=-19.90\pm 0.29$ mag, roughly half a magnitude brighter than typical SNe Ia of similar color-stretch values.
@@
Filter JD peak Peak app. Mag. $\Delta m_{15}$ (mag)
@@ 4.1.2 Estimating the time of first light, $t_{first}$
This pre-fireball emission should not be confused with the early flux excess seen in many early SNe Ia light curves, as the two phenomena are characteristically distinct. The early flux excess is typically attributed to the interaction of the SN ejecta with a non-degenerate companion star (2010ApJ...708.1025K), or with circumstellar material (e.g., Piro2016; Maeda2023; 2023MNRAS.522.6035M; 2018ApJ...864L..35S). It is estimated that approximately $20-30\%$ of all SNe Ia exhibit such early-excess features (2022MNRAS.513.3035M; 2022MNRAS.512.1317D). However, in certain subclasses of SNe Ia – such as the peculiar 2002es-like, 2003fg-like, and luminous 1991T-like events – the occurrence rate of early-excess emission appears to be considerably higher (2018ApJ...865..149J; 2024ApJ...966..139H). Crucially, unlike these early-excess signatures, the pre-fireball emission observed in SN 2021hem likely arises from a distinct physical mechanism, as discussed earlier. An early flux excess due to interaction typically emerges within a few days of explosion and appears superimposed on the power-law rise of the early light curve. This is because the short-lived excess flux from the interaction and the underlying “fireball” emission are both consequences of the expanding ejecta. In other words, the interaction can not occur without the “fireball” expansion. In contrast, the early emission seen in SN 2021hem occurs even before the onset of the fireball’s power-law flux rise, clearly distinguishing it from the early excesses or bumps arising from interaction in other SNe Ia. Moreover, 2017Natur.550...80J argued, based on their models, that an early excess arising from interaction would inevitably show a blue color evolution, as was observed in the $(\mbox{$B\!-\!V$})_{0}$ color of SN iPTF14atg. 2020ApJ...902...48B has also shown using ejecta-companion interaction models that the $(g-r)_{0}$ colors at the time of interaction is expected to be ${\sim}-0.5$ mag. In the case of SN 2021hem the $(g-r)_{0}$ color of $0.16\pm 0.33$ mag, computed from first epoch of $g$ and $r$-band light curves, does not show any blue excess. This color is broadly consistent with, or perhaps marginally redder than, the early $(g-r)_{0}$ colors typically observed in normal SNe Ia (2020ApJ...902...48B). Therefore, considering all these factors, the interaction scenario is disfavored for the pre-fireball emission seen in SN 2021hem.
@@
Regardless of the origin, this early activity, which can be described as early flux excess or pre-fireball flux emission as in SN 2021hem, appears to be a ubiquitous feature among 2003fg-like SNe. In addition to SN 2021hem, other 2003fg-like supernovae with early-time observations and adequate photometric cadence – SNe 2020hvf, 2021zny, 2022ilv, and 2021qvo (2021ApJ...923L...8J; 2024ApJ...966..139H; 2025arXiv250813263A) – have all exhibited an early bump in their light curves.
@@
Figure 5: UVOIR bolometric light curve of SN 2021hem fit with a radioactive 56Ni decay model. The time of explosion is taken as the epoch of first detection, which is $\approx 17.9$ days before the $B$-band maximum. Only data beyond 60 days are used for fitting the model. The UVOIR light curve of SN 2021hem (top panel), its time-weighted integrated luminosity (middle panel), and the ratio $t^{2}L/\left(\int Lt\penalty 10000\ dt\right)$ (bottom panel), which is defined independent of the 56Ni mass are shown.